# When to use Graphs in RAG: A Comprehensive Analysis for Graph Retrieval-Augmented Generation

### v1

@@ A.2 Why is it important to include tasks with varying complexity?
A.3 Why construct two datasets from both novels and the medical corpus?
@@
Including two distinct datasets, medical guidelines and unstructured novels, is essential to evaluate GraphRAG models under conditions that mirror real-world knowledge ecosystems. Medical corpora provide explicit, hierarchical relationships, testing a model’s ability to navigate rigid domain logic and standardized protocols. Conversely, novels introduce implicit, context-dependent dependencies, like socio-historical factors shaping character decisions, challenging models to infer latent connections without predefined rules. This approach ensures the benchmark assesses both precision in following formal hierarchies and adaptability in interpreting ambiguous, open-ended contexts, critical for applications where models must integrate structured expertise with unstructured, real-world narratives.
A.4 What is the advantage of using stage-specific evaluation metrics?
@@
Stage-specific evaluation metrics are crucial because they provide granular insights into how well a GraphRAG model performs at each phase of its pipeline, rather than relying solely on end-to-end output metrics like answer accuracy. Traditional benchmarks often treat the entire process as a “black box,“ obscuring whether failures stem from flawed knowledge graph construction, suboptimal retrieval, or weak reasoning. By designing metrics tailored to individual stages, such as graph completeness during logic mining, retrieval relevance in evidence collection, or contextual coherence in question generation, we isolate and diagnose weaknesses in specific components.
A.5 Comparison with existing benchmarks and analysis papers.
@@
Existing studies [22, 65] focus on architectural comparisons using homogeneous datasets, missing how models synthesize hierarchical expertise and unstructured narratives. To this end, we propose GraphRAG-Bench, a comprehensive benchmark designed to evaluate GraphRAG models on deep reasoning. It features hybrid corpora (medical guidelines + novels) with tasks of increasing complexity and stage-specific metrics to expose why models fail, whether in graph construction, knowledge retrieval, or contextual synthesis. Leveraging this novel benchmark, we systematically investigate the conditions when GraphRAG surpasses traditional RAG systems and the underlying reasons for its success, offering guidelines for its practical application.
Appendix B Benchmark Construction
@@
To address the critical limitations of existing RAG evaluation frameworks, we present a novel benchmark to assess GraphRAG systems through comprehensive task hierarchies and structured knowledge integration. Specifically, our benchmark is constructed through six stages that systematically integrate domain-specific logical hierarchies and contextual dependencies to enable precise control over question difficulty. The example and related prompts are illustrated in Figures 10 and 13.
B.1 Corpus Collection
@@ B.2 Logic Mining
GraphRAG (local) [13] 49.29 26.11 50.93 24.09 64.40 75.58 39.10 55.44 35.65
GraphRAG (global) [13] 36.92 17.32 43.17 15.12 56.87 80.55 41.11 75.15 30.34
HippoRAG [20] 52.93 26.65 38.52 11.16 48.70 85.55 38.85 71.53 38.97
HippoRAG2 [21] 60.14 31.35 53.38 33.42 64.10 70.84 48.28 49.84 30.95
LightRAG [19] 58.62 35.72 49.07 24.16 48.85 63.05 23.80 57.28 25.01
Fast-GraphRAG [11] 56.95 35.90 48.55 21.12 56.41 80.82 46.18 57.19 36.99
Graph RAG RAPTOR [48] 49.25 23.74 38.59 11.66 47.10 82.33 38.01 70.85 35.88
@@
GraphRAG (local) [13] 38.63 26.80 47.04 21.99 41.87 22.98 53.11 32.65 39.42
GraphRAG (global) [13] 16.42 46.00 15.61 52.75 - - - - -
HippoRAG [20] 56.14 20.95 55.87 13.57 59.86 62.73 64.43 69.21 65.56
HippoRAG2 [21] 66.28 36.69 61.98 36.97 63.08 46.13 68.05 58.78 51.54
LightRAG [19] 63.32 37.19 61.32 24.98 63.14 51.16 67.91 78.76 51.58
Fast-GraphRAG [11] 60.93 31.04 61.73 21.37 67.88 52.07 65.93 56.07 44.73
GraphRAG RAPTOR [48] 54.07 17.93 53.20 11.73 58.73 78.28 62.38 58.98 63.63
@@
GraphRAG (local) [13] 39.89 25.93 46.12 27.70 65.28 69.10 39.57 54.64 32.42
GraphRAG (global) [13] 39.89 11.54 42.22 19.10 60.41 75.45 36.59 84.60 34.30
HippoRAG2 [21] 54.79 30.16 50.45 30.30 61.14 68.99 40.52 52.24 32.05
LightRAG [19] 44.00 12.22 40.27 9.91 52.07 86.20 39.74 78.73 39.67
Fast-GraphRAG [11] 60.08 41.31 53.81 30.43 62.82 74.45 47.60 57.99 30.31
Graph RAG RAPTOR [48] 41.12 11.91 41.15 9.58 52.03 81.04 38.56 63.10 34.77
@@
GraphRAG (local) [13] 49.24 29.17 61.64 27.40 59.01 37.12 61.70 33.39 42.24
GraphRAG (global) [13] 40.04 17.01 61.40 21.08 51.10 38.69 59.20 - 43.17
HippoRAG2 [21] 64.50 33.92 64.05 33.02 64.71 48.27 60.77 39.33 36.76
LightRAG [19] 64.43 40.37 64.66 28.26 69.37 59.81 63.25 70.84 45.12
Fast-GraphRAG [11] 62.03 43.59 62.40 29.42 65.09 46.29 62.98 31.98 35.83
GraphRAG RAPTOR [48] 50.48 13.49 52.86 12.96 58.94 78.63 61.45 49.91 54.46
Table 6: Retrieval performance across tasks with varying complexity from different datasets.
Novel Dataset
RAG (w/o rerank) 61.37 74.66 59.80 80.82 69.08 80.05 32.48 82.84
GraphRAG (local)[13] 61.04 27.30 73.03 39.09 82.02 43.13 53.55 35.07
GraphRAG (global)[13] 42.27 9.37 86.68 14.36 89.69 15.35 83.14 19.40
HippoRAG[20] 80.44 56.34 87.91 58.75 90.95 59.46 65.51 46.64
HippoRAG2[21] 70.29 79.25 69.77 85.75 82.50 87.82 42.18 79.10
LightRAG[19] 73.69 33.08 85.52 37.46 87.59 38.02 71.72 38.06
Fast-GraphRAG[11] 64.48 47.86 73.51 55.21 78.58 49.74 56.31 46.27
Graph RAG RAPTOR[48] 62.14 54.08 67.80 61.26 75.79 63.00 58.66 58.46
Medical Dataset
RAG (w/o rerank) 86.24 63.71 84.97 84.11 84.14 89.94 44.88 58.73
GraphRAG (local)[13] 38.06 5.67 61.32 4.25 59.66 5.24 66.59 2.76
GraphRAG (global)[13] 65.98 7.46 78.46 11.72 - - - -
HippoRAG [20] 87.25 52.44 83.80 42.19 83.46 49.13 81.66 45.03
HippoRAG2 [21] 78.70 87.96 77.00 80.94 77.40 86.85 61.12 78.64
LightRAG [19] 80.32 41.27 82.91 42.79 85.71 43.11 81.34 45.17
Fast-GraphRAG [11] 66.82 45.86 74.93 38.80 77.27 47.58 62.99 25.15
Graph RAG RAPTOR [48] 85.40 69.38 89.70 53.20 88.86 58.73 72.70 52.71
B.3 Evidence Collection
B.4 Question Generation
B.5 Check & Correct
B.6 Refinement
Appendix C Dataset Statistics and Visualization
@@
Benchmark Avg. Entities Avg. Relations Prop. of Non-isolated Entities Avg. Degree Prop. Degree > 1 Prop. Degree > 2 Prop. Degree > 3 Avg. Component Size
@@
Questions statistics We first categorize questions for existing benchmarks based on a predefined taxonomy of difficulty levels, with classification results shown in Figure 12. Our analysis reveals a significant imbalance in the distribution of question levels across current benchmarks. Specifically, UltraDomain focuses mostly on Contextual Summarization questions (97%), while HotpotQA mainly contains Fact Retrieval questions (78.2%), lacking coverage of deeper logical reasoning tasks. Although MultiHop-RAG balances Complex Reasoning and Contextual Summarization questions better, it entirely lacks basic Fact Retrieval questions, making its evaluation coverage incomplete.
@@
Appendix D Evaluation Metrics

### v2

@@ A.2 Why is it important to include tasks with varying complexity?
A.3 How to control the complexity of each task?
GraphRAG-Bench leverages structural information to control complexity. We begin by extracting information from corpora to build ontologies or knowledge graphs. Taking the Novel dataset as an example, we first construct a knowledge graph from the corpus, then form logic chains using relevant triples from this graph. Then, we control difficulty by adjusting the number of involved triples and the information span they cover. Specifically, we define question complexity through two dimensions: Knowledge Breadth (measured by the count of triples required to answer a question) and Reasoning Depth (measured by the number of inference hops between these triples). The relevant statistical results are shown in Table 4.
Table 4: Problem Complexity Statistics of GraphRAG-Bench.
Problem Complexity Fact Retrieval Complex Reasoning Contextual Summarize Creative Generation Novel Dataset    Knowledge Breadth 1.40 2.60 3.51 7.11    Reasoning Depth 1.69 6.25 4.64 7.81 Medical Dataset    Knowledge Breadth 1.25 3.45 5.1 10.14    Reasoning Depth 1.82 5.23 4.27 8.27
A.4 Why construct two datasets from both novels and the medical corpus?
@@
Including two distinct datasets, medical guidelines and unstructured novels, is essential to evaluate GraphRAG models under conditions that mirror real-world knowledge ecosystems. Medical corpora provide explicit, hierarchical relationships, testing a model’s ability to navigate rigid domain logic and standardized protocols. Conversely, Novel corpora introduce implicit, context-dependent dependencies, like socio-historical factors shaping character decisions, challenging models to infer latent connections without predefined rules. This approach ensures the benchmark assesses both precision in following formal hierarchies and adaptability in interpreting ambiguous, open-ended contexts, critical for applications where models must integrate structured expertise with unstructured, real-world narratives.
A.5 Why Not Use Other Formats for Question Construction?
In this paper, we aim to figure out in which scenarios do graph structures provide measurable benefits for RAG systems. In other words, we care more about the retrieval and reasoning complexity of task, instead of the task type or question format. It is because the specific task type or format (like QA, multiple-choice, or fact-checking) doesn’t really change: 1) reasoning difficulty: the reasoning steps the model needs to take, 2) retrieval difficulty: how to locate scattered information from the corpus.
Given a question of “What famous universities are in San Diego?”. Whether this is asked as: An open QA question, or A multiple-choice question (e.g., “Is UCSD in San Diego? A) Yes B) No”), or A fact-checking task (e.g., “Check this: UCSD has a campus in San Diego”)… doesn’t change the core need: The RAG system must still find information about San Diego and reason to answer correctly.
A.6 What is the advantage of using stage-specific evaluation metrics?
@@
Stage-specific evaluation metrics are crucial because they provide granular insights into how well a GraphRAG model performs at each phase of its pipeline, rather than relying solely on end-to-end output metrics like answer accuracy. Traditional benchmarks often treat the entire process as a “black box”, obscuring whether failures stem from flawed knowledge graph construction, suboptimal retrieval, or weak reasoning. By designing metrics tailored to individual stages, such as graph completeness during logic mining, retrieval relevance in evidence collection, or contextual coherence in question generation, we isolate and diagnose weaknesses in specific components.
A.7 Comparison with existing benchmarks and analysis papers.
@@
Existing studies (Han et al., 2025; Zhou et al., 2025) focus on architectural comparisons using homogeneous datasets, missing how models synthesize hierarchical expertise and unstructured narratives. To this end, we propose GraphRAG-Bench, a comprehensive benchmark designed to evaluate GraphRAG models on deep reasoning. It features hybrid corpora (medical guidelines + novels) with tasks of increasing complexity and stage-specific metrics to expose why models fail, whether in graph construction, knowledge retrieval, or contextual synthesis. Leveraging this novel benchmark, we systematically investigate the conditions when GraphRAG surpasses traditional RAG systems and the underlying reasons for its success, offering guidelines for its practical application.
Appendix B Takeaway Findings
In this paper, we not only build GraphRAG-Bench to evaluate existing GraphRAG systems, but more importantly, we provide insightful recommendations for future GraphRAG research, as illustrated in Figure 11.
Prioritize precise retrieval: Effective frameworks should focus on how to maximize key information retrieval while minimizing redundancy. It’s critical to pinpoint the key facts needed to answer a question, while at the same time avoiding pulling in unnecessary details. This keeps the context clean and focused, which helps the model reason better and improves overall efficiency.
Build quality graphs, not just large ones: While GraphRAG forms knowledge graphs (entities/relationships) for efficient searching, more relationships $\neq$ better performance. Optimal graphs require tightly connected communities, which create denser structures rich in implicit multi-hop knowledge – enabling faster graph traversal.
Actively manage context growth: Unlike traditional RAG (fixed context via vector search), GraphRAG retrieves entities, relationships, and raw text snippets, risking sudden context explosion and hig reasoning costs. Future solutions need search boundaries to curb context growth and significantly lower costs.
Figure 11: An overview of building an effective GraphRAG system. It consists of two key parts: (i) the crucial principles for system design (left), and (ii) the most suitable application scenarios (right).
Appendix C Benchmark Construction
@@
To address the critical limitations of existing RAG evaluation frameworks, we present a novel benchmark to assess GraphRAG systems through comprehensive task hierarchies and structured knowledge integration. Specifically, our benchmark is constructed through six stages that systematically integrate domain-specific logical hierarchies and contextual dependencies to enable precise control over question difficulty. The example and related prompts are illustrated in Figures 12 and 15.
C.1 Corpus Collection
@@ C.2 Logic Mining
MS-GraphRAG (local) (Edge et al., 2024) 49.29 26.11 50.93 24.09 64.40 75.58 39.10 55.44 35.65
MS-GraphRAG (global) (Edge et al., 2024) 36.92 17.32 43.17 15.12 56.87 80.55 41.11 75.15 30.34
HippoRAG (Gutiérrez et al., 2024) 52.93 26.65 38.52 11.16 48.70 85.55 38.85 71.53 38.97
HippoRAG2 (Gutiérrez et al., 2025) 60.14 31.35 53.38 33.42 64.10 70.84 48.28 49.84 30.95
LightRAG (Guo et al., 2024) 58.62 35.72 49.07 24.16 48.85 63.05 23.80 57.28 25.01
Fast-GraphRAG (CircleMind-AI, 2024) 56.95 35.90 48.55 21.12 56.41 80.82 46.18 57.19 36.99
Graph RAG RAPTOR (Sarthi et al., 2024) 49.25 23.74 38.59 11.66 47.10 82.33 38.01 70.85 35.88
Lazy-GraphRAG (Darren Edge, 2024) 51.65 36.97 49.22 23.48 58.29 76.94 43.23 50.69 39.74
KGP (Wang et al., 2024b) 54.15 24.73 46.31 16.91 51.21 64.34 40.37 52.55 34.65
StructRAG (Li et al., 2024) 53.84 26.73 46.27 23.49 54.28 63.56 42.16 52.68 36.75
KET-RAG (Huang et al., 2025) 55.39 27.39 36.59 25.98 52.47 69.24 46.03 36.72 33.68
@@
MS-GraphRAG (local) (Edge et al., 2024) 38.63 26.80 47.04 21.99 41.87 22.98 53.11 32.65 39.42
MS-GraphRAG (global) (Edge et al., 2024) 16.42 46.00 15.61 52.75 19.82 - 20.81 - 13.64
HippoRAG (Gutiérrez et al., 2024) 56.14 20.95 55.87 13.57 59.86 62.73 64.43 69.21 65.56
HippoRAG2 (Gutiérrez et al., 2025) 66.28 36.69 61.98 36.97 63.08 46.13 68.05 58.78 51.54
LightRAG (Guo et al., 2024) 63.32 37.19 61.32 24.98 63.14 51.16 67.91 78.76 51.58
Fast-GraphRAG (CircleMind-AI, 2024) 60.93 31.04 61.73 21.37 67.88 52.07 65.93 56.07 44.73
GraphRAG RAPTOR (Sarthi et al., 2024) 54.07 17.93 53.20 11.73 58.73 78.28 62.38 59.98 63.63
Lazy-GraphRAG (Darren Edge, 2024) 60.25 31.66 47.82 22.68 57.28 55.92 62.22 30.95 43.79
KGP (Wang et al., 2024b) 52.34 21.34 51.53 11.69 54.51 62.40 63.77 45.25 35.55
StructRAG (Li et al., 2024) 55.38 27.53 56.17 22.79 62.48 65.66 60.21 42.35 45.76
KET-RAG (Huang et al., 2025) 60.35 31.99 39.56 19.52 45.27 29.04 43.04 33.67 31.93
Table 6: Results of Retrieval performance using GPT-4o-mini, covering tasks of varying complexity.
Novel Dataset
RAG (w/o rerank) 61.37 74.66 59.80 80.82 69.08 80.05 32.48 82.84
MS-GraphRAG (local)(Edge et al., 2024) 61.04 27.30 73.03 39.09 82.02 43.13 53.55 35.07
MS-GraphRAG (global)(Edge et al., 2024) 42.27 09.37 86.68 14.36 89.69 15.35 83.14 19.40
HippoRAG(Gutiérrez et al., 2024) 80.44 56.34 87.91 58.75 90.95 59.46 65.51 46.64
HippoRAG2(Gutiérrez et al., 2025) 70.29 79.25 69.77 85.75 82.50 87.82 42.18 79.10
LightRAG(Guo et al., 2024) 73.69 33.08 85.52 37.46 87.59 38.02 71.72 38.06
Fast-GraphRAG(CircleMind-AI, 2024) 64.48 47.86 73.51 55.21 78.58 49.74 56.31 46.27
Graph RAG RAPTOR(Sarthi et al., 2024) 62.14 54.08 67.80 61.26 75.79 63.00 58.66 58.46
Lazy-GraphRAG (Darren Edge, 2024) 59.25 30.76 57.73 42.98 77.38 43.62 55.24 31.94
KGP (Wang et al., 2024b) 55.71 23.71 63.51 31.96 61.54 64.20 67.57 35.52
StructRAG (Li et al., 2024) 55.38 27.53 56.17 34.79 62.48 65.66 60.21 42.35
KET-RAG (Huang et al., 2025) 63.55 39.11 56.93 32.59 67.35 39.05 53.40 36.74
Medical Dataset
RAG (w/o rerank) 86.24 63.71 84.97 84.11 84.14 89.94 44.88 58.73
MS-GraphRAG (local)(Edge et al., 2024) 38.06 05.67 61.32 04.25 59.66 05.24 66.59 02.76
MS-GraphRAG (global)(Edge et al., 2024) 65.98 07.46 78.46 11.72 89.06 11.72 85.28 02.73
HippoRAG (Gutiérrez et al., 2024) 87.25 52.44 83.80 42.19 83.46 49.13 81.66 45.03
HippoRAG2 (Gutiérrez et al., 2025) 78.70 87.96 77.00 80.94 77.40 86.85 61.12 78.64
LightRAG (Guo et al., 2024) 80.32 41.27 82.91 42.79 85.71 43.11 81.34 45.17
Fast-GraphRAG (CircleMind-AI, 2024) 66.82 45.86 74.93 38.80 77.27 47.58 62.99 25.15
Graph RAG RAPTOR (Sarthi et al., 2024) 85.40 69.38 89.70 53.20 88.86 58.73 72.70 52.71
Lazy-GraphRAG (Darren Edge, 2024) 74.29 19.90 78.65 17.50 78.72 21.35 83.41 15.09
KGP (Wang et al., 2024b) 57.51 27.34 53.51 26.59 59.38 56.20 68.42 43.85
StructRAG (Li et al., 2024) 63.25 37.26 61.75 35.68 62.55 32.01 62.76 46.75
KET-RAG (Huang et al., 2025) 86.44 57.07 80.62 30.86 89.07 44.59 44.06 32.38
C.3 Evidence Collection
C.4 Question Generation
C.5 Check & Correct
C.6 Refinement
@@
MS-GraphRAG (local) (Edge et al., 2024) 39.89 25.93 46.12 27.70 65.28 69.10 39.57 54.64 32.42
MS-GraphRAG (global) (Edge et al., 2024) 39.89 11.54 42.22 19.10 60.41 75.45 36.59 84.60 34.30
HippoRAG2 (Gutiérrez et al., 2025) 54.79 30.16 50.45 30.30 61.14 68.99 40.52 52.24 32.05
LightRAG (Guo et al., 2024) 44.00 12.22 40.27 9.91 52.07 86.20 39.74 78.73 39.67
Fast-GraphRAG (CircleMind-AI, 2024) 60.08 41.31 53.81 30.43 62.82 74.45 47.60 57.99 30.31
Graph RAG RAPTOR (Sarthi et al., 2024) 41.12 11.91 41.15 9.58 52.03 81.04 38.56 63.10 34.77
@@
MS-GraphRAG (local) (Edge et al., 2024) 49.24 29.17 61.64 27.40 59.01 37.12 61.70 33.39 42.24
MS-GraphRAG (global) (Edge et al., 2024) 40.04 17.01 61.40 21.08 51.10 38.69 59.20 - 43.17
HippoRAG2 (Gutiérrez et al., 2025) 64.50 33.92 64.05 33.02 64.71 48.27 60.77 39.33 36.76
LightRAG (Guo et al., 2024) 64.43 40.37 64.66 28.26 69.37 59.81 63.25 70.84 45.12
Fast-GraphRAG (CircleMind-AI, 2024) 62.03 43.59 62.40 29.42 65.09 46.29 62.98 31.98 35.83
GraphRAG RAPTOR (Sarthi et al., 2024) 50.48 13.49 52.86 12.96 58.94 78.63 61.45 49.91 54.46
*Due to time and resource constraints, we tested our benchmark on a representative set of GraphRAG models.
Appendix D Extending GraphRAG-Bench to New Domains
In this paper, we used a medical dataset to represent domain knowledge in the initial version of GraphRAG-Bench. The method we constructed the data can be used to extend the benchmark into other fields, such as law and finance. For researchers who wish to add other domains in future work, we provide some suggested methods.
For the legal domain, we offer the following suggestions: 1) Corpus Selection: We suggest collecting data from the following sources: a) EU Case Law which contains 29.8K EU court decisions, mainly from the Court of Justice (CJEU), published in EUR-Lex. b) UK Case Law which contains 47K UK court decisions from the British and Irish Legal Information Institute (BAILII) database. c) US Case Law which contains 4.6M US decisions (opinions) from Court Listener, a web database hosted by the Free Law Project. 2) Data Mining Method: We recommend a method that integrates both ontology and logic chains. Ontology provides a structured representation of legal regulations and judicial interpretations by defining entities, actions, relationships, and the conditions under which specific laws apply. Logic chains, in parallel, capture the legal reasoning process and model potential multi-hop relationships between different laws.
For the finance domain, we suggest the following: 1) Corpus Selection: We recommend focusing on publicly traded companies from the S&P 500 list for the period between 2015 and 2024. For these companies, we can collect their annual reports, quarterly reports, and reports on unscheduled events from the same period. These documents are primarily available in the EDGAR database. 2) Data Mining Method: We suggest using an ontology to map company information, such as business structures and financial metrics, to a unified schema. It is worth noting that the financial domain contains a large amount of numerical data, which may require further processing. The details statistics of Legal and Financial Corpora are in Table 8.
Table 8: Statistics of Legal and Financial Corpora.
Source Numbers Tokens
EU Case Law 14.9K 89.25M
UK Case Law 11.7K 92.10M
U.S. Case Law 46.1K 114.23M
Total 72.7K 295.58M
(a) Statistics of Legal Corpora.
Report Type Numbers Tokens
Annual Report 192 18.72M
Quarterly Report 588 22.93M
Current Report 1427 74.21M
Total 2207 115.86M
(b) Statistics of Financial Corpora.
Appendix E Dataset Statistics and Visualization
@@
Benchmark
Avg.
Entities
Avg.
Relations
Prop. of
Non-isolated Entities
Avg.
Degree
Prop.
Degree > 1
Prop.
Degree > 2
Prop.
Degree > 3
Avg.
Component Size
@@
Questions statistics We first categorize questions for existing benchmarks based on a predefined taxonomy of difficulty levels, with classification results shown in Figure 14. Our analysis reveals a significant imbalance in the distribution of question levels across current benchmarks. Specifically, UltraDomain focuses mostly on Contextual Summarize questions (97%), while HotpotQA mainly contains Fact Retrieval questions (78.2%), lacking coverage of deeper logical reasoning tasks. Although MultiHop-RAG balances Complex Reasoning and Contextual Summarize questions better, it entirely lacks basic Fact Retrieval questions, making its evaluation coverage incomplete.
@@
Appendix F Evaluation Metrics
Retrieval Performance
To evaluate the retrieval performance of GraphRAG, we argue that an effective system should both guarantee the completeness of retrieved information and reduce irrelevant content. We introduce two corresponding retrieval-quality-based metrics as detailed below
$\textsc{Context Relevance}=\frac{1}{|\mathcal{C}|}\sum_{c\in\mathcal{C}}\operatorname{R}(c,Q,\mathcal{E}),$ (3)
where $\mathcal{C}$ denotes the set of retrieved contexts, $Q$ represents the question, $\mathcal{E}$ denotes the set of evidence, and the operator $\operatorname{R}(\cdot)$ determines whether a context $c$ is relevant to the question $Q$ and the evidence $\mathcal{E}$.
$\textsc{Evidence Recall}=\frac{1}{|\mathcal{R}|}\sum_{c\in\mathcal{R}}\mathds{1}\left(\operatorname{S}\left(c,\mathcal{C}\right)\right),$ (4)